A father said to his son, " I was twice as old as you are now on the day that you were born. You will be just what i was then when fourteen years are gone ".
What are the present ages of the father and son.

Answer by ankor@dixie-net.com(22740)

A father said to his son, " I was twice as old as you are now on the day that you were born.
You will be just what i was then when fourteen years are gone ".
What are the present ages of the father and son.
:
Let s = son's present age
Father was twice as old when son was born therefore
2s + s = 3s; father's present age
:
"You will be just what i was then when fourteen years are gone ".
2s = s + 14
2s - s = 14
s = 14 yrs is son's present age
then
3(14) = 42 yrs is father's present age